OpenStudy (anonymous):

A Honda moving at 100km/hr is 80 km behind a Ford moving in the same direction at 60 km/hr. How far does the Honda travel before it catches up with the Ford?

OpenStudy (anonymous):

Let us fix an x-axis: in the origin of this axis there is the Honda, at the position of 80 km there is the Ford. The two laws of motion are: \[x(H)= 100 km/h *t\] \[x(F)= 60 km/h *t + 80 km\] the Honda will catch the ford when the position is the same: solve the equation \[60 km/h *t + 80 km= 100 km/h *t\] then obtain t=2 h substitute in x(H) and obtain that the Honda has travelled 50 km before catching the Ford!

OpenStudy (stormfire1):

I'd make this problem a lot simpler and use relative velocities. The relative velocity between the cars is 100km/h-60km/h = 40km/h. This is the amount the Honda is gaining on the Ford. Since the Ford is 80km away, it will take 2 hours at 40km/h.